OpenStudy (anonymous):

here is a nice question that i'm working on right now : find derivative of : \[f(x) = \int\limits_{x}^{2x}\sin(t^2)dt \]

OpenStudy (anonymous):

it's pretty easy. you should use the following formula - \[\int\limits_{\alpha}^{\beta} f(x) = f(\beta) f'(\beta) - f(\alpha)f'(\alpha)\] So your answer should be - \[2\sin (4x^2) - \sin(x^2)\]

OpenStudy (anonymous):

x appaers in both limits so we can make that : \[\int\limits_{x}^{2x}\sin(t^2)dt =\int\limits_{x}^{0}\sin(t^2)dt + \int\limits_{0}^{2x}\sin(t^2)dt \]

OpenStudy (anonymous):

Yup^^Leibniz One

OpenStudy (anonymous):

\[d/dx(-\int\limits_{0}^{x}\sin(t^2)dt) = -\sin(x^2)\]

OpenStudy (anonymous):

u = 2x \[y = \int\limits_{0}^{u}\sin(t^2)dt\] so that \[dy/dx = (dydu) / (dudx) = ( \sin(u^2)) (2) = 2\sin(4x^2) \]

OpenStudy (anonymous):

yup same thing Lfc :D: D

OpenStudy (anonymous):

2 sin(4x2) - sin(x2)